OpenStudy (anonymous):

Which of the following equals 3x^2 - 10x - 8 when factored completely? can you show it step by step?

OpenStudy (anonymous):

x = -b + SQRT(b^2 - 4ac) /2a , -b + SQRT(b^2 - 4ac) /2a = (10 + SQRT (100 +96)) /6 , (10 - SQRT (100 +96)) /6 = 4,-4/6 (x-4) (x+4/6)